# Multi-Modal Graph Convolutional Network with Sinusoidal Encoding for Robust Human Action Segmentation

### v1

@@ I INTRODUCTION
Our preliminary studies [1] on human-robot collaboration reveal that encoder-decoder graph convolutional networks excel at parsing spatiotemporal information by representing human skeleton and object center points as graph nodes. Through attention mechanisms, these networks dynamically update the relationships between nodes, capturing the evolving interactions between human and objects. However, joint misdetections and tracking errors in skeleton data often lead to over-segmentation, fragmenting continuous actions into disjointed segments and disrupting the temporal coherence of action sequences. This limitation becomes particularly critical in human-object interaction scenarios, where both human pose estimation (skeleton) and object detections (bounding boxes) introduce noise. Object misdetections, such as missing and inaccurately localized bounding boxes, lead to inconsistencies in action segmentation.
@@
Recent advances in multi-modal networks [2, 3] have demonstrated promise in addressing noise-induced over-segmentation by integrating structured 3D motion data (e.g., skeletal trajectories and object centroids) with pixel-based visual streams. However, integrating pixel-based RGB features with 3D position information presents significant challenges. The two modalities exhibit divergent characteristics: 3D position data are spatially sparse and structured, while image features are spatially dense and unstructured. Directly concatenating or aligning these representations leads to feature incompatibility. To bridge this gap, we introduce a sinusoidal joint encoding strategy that projects 3D coordinates into a continuous sin-cos space. This transformation downscales noisy joint trajectories while preserving spatial relationships, enabling seamless fusion with visual features.
@@
A second challenge arises from computational efficiency: processing full-frame video data at high temporal resolutions is prohibitively expensive for real-time systems, while naively fusing features across different temporal resolutions degrades performance. A common approach is to upsample low-frequency image features to match the motion sequence’s time scale and merge them using a late fusion strategy [3]. However, this final fusion step often results in feature loss due to the mismatch in the fineness of features, and it overlooks fine-grained temporal dependencies and misaligns crucial visual and motion cues. To address this, we propose a Multi-Modal Graph Convolutional Network (MMGCN) that processes high-frequency motion data via a hierarchical graph convolutional network and strategically integrates high-frequency motion features into low-frequency (1 fps) visual cues at a middle stage, ensuring a more coherent fusion with processed motion information later in the pipeline. By processing the two modalities in different temporal resolutions, our method effectively preserves fine-grained temporal dependencies and aligns visual motion cues efficiently.
@@ II-A Human Action Segmentation
Human Action Segmentation, which aims to temporally localize and classify sub-actions within continuous sequences, has evolved from early template-based methods [4, 5] to modern image-based learning architectures [6, 7]. Temporal Convolutional Networks (TCNs) [6] pioneered frame-wise prediction using dilated convolutions, while MS-TCN [7] extended this with multi-stage refinement. Recent works like Action Segment Refinement Framework [8] combine boundary detection and classification to improve temporal localization. Despite progress, these methods suffer from high computational costs, sensitivity to background clutter, and the inclusion of redundant visual information, making them inefficient for real-time applications.
To address these challenges, graph-based networks are considered as a more efficient alternative by modeling human motion as structured spatial and temporal graphs, significantly reducing background noise and computational overhead. Early works, such as Spatial-Temporal Graph Convolutional Network (ST-GCN) [9], introduced graph convolutional networks to sequentially capture joint-level motion patterns across both spatial and temporal dimensions. Subsequent improvements, like Two-Stream Adaptive Graph Convolutional Network (2s-AGCN) [10] and Multi-Scale Graph Network (MS-G3D) [11], incorporated attention mechanisms and multi-scale feature extraction to enhance action recognition. More recently, methods such as Pyramid Graph Convolutional Network (PGCN) [12] and Temporal Fusion Graph Convolutional Network (TFGCN) [1] have explored temporal modeling techniques to further improve segmentation accuracy.
@@ II-B Multi-Modal Models
Multi-modal learning is a promising method for action recognition, leveraging complementary strengths of visual (RGB/depth) and skeletal modalities to overcome individual limitations. Early works, like [13, 14] fuse images/depths and skeletons per frame to improve the recognition results. Recent advancements focus on intermediate fusion strategies to better align cross-modal features. Kim et al. [15] proposed a cross-modal transformer that attends to spatial correspondences between skeleton joints and image regions, improving recognition consistency by dynamically weighting discriminative features. Similarly, Wang et al. [16] integrates skeleton joints with RGB patches using a late fusion technique, but its reliance on dense pixel processing limits scalability. To address efficiency, PoseC3D [2] encodes skeletons as pseudo-3D heatmaps and fuses them with RGB features via 3D convolutions, achieving a balance between accuracy and speed. More recently, Tani Hiroaki [3] introduced a temporal attention mechanism that selects one representative image for the entire motion sequence capturing the minimal appearance features necessary for action recognition. While this approach is effective for short-term action recognition, it becomes computationally intensive for long-duration actions.
@@ II-C Data Augmentation
Data augmentation plays a pivotal role in improving the generalization and robustness of temporal models, particularly for action recognition and segmentation tasks. Traditional techniques, such as spatial augmentations (e.g., cropping, flipping) [17] and temporal perturbations (e.g., frame skipping, time warping) [18], focus on enhancing spatial invariance or simulating temporal variations. While effective for classification, these methods often disrupt the temporal coherence required for precise segmentation, where smooth transitions between sub-actions are critical.
Recent approaches adapt image-domain augmentation strategies to temporal tasks. Mixup [19] and CutMix [20], which linearly interpolate inputs and labels, have been extended to videos by blending random clips (e.g., VideoMix [21]). However, these methods prioritize spatial or short-term temporal invariance, inadvertently introducing abrupt label transitions that exacerbate over-segmentation artifacts. For instance, TemporalMix [22] downsamples two segments by half, interpolates the larger segment to match the smaller one to preserve the natural smoothness of human actions, and then combines them into a new sequence. However, its downsampling process disrupts the natural tempo of human actions, leading to inconsistent predictions in speed-sensitive tasks.
@@ III Multi-Modal Graph Convolutional Network
The idea of the Multi-Modal Graph Convolutional Network is jointly optimizing dense spatial features from visual data and dense temporal dynamics from the motion sequences (skeleton and object). Motion inputs (e.g., 30 fps) provide fine-grained motion patterns but suffer from joint tracking noise, while sparse RGB frames (1 fps) offer stable spatial context but lack temporal granularity. Our model bridges these modalities through two core mechanisms: a sinusoidal encoder and temporal feature refinement.
@@ III-A Sinusoidal Encoder
To mitigate spatial jitter and enhance compatibility with visual features, we transform the joint position information in the camera coordinate into a continuous sin-cons representation. Given a 3D joint $P_{i}=(X_{i},Y_{i},Z_{i})$, where $P_{i}\in\mathbb{R}^{1x3}$, the encoder maps each coordinate into a high-dimensional space using sinusoidal and cosine functions. The encoder first scales these coordinates by a frequency factor determined by the parameters $\alpha$ and $\beta$. It then computes sinusoidal $sin\_e$ and cosine $con\_e$ embeddings for each dimension, resulting in a high-dimensional vector that captures both the position and its spatial context. Mathematically, for each coordinate $c\in\{X_{i},Y_{i},Z_{i}\}$, the encoder computes:
@@ III-B Temporal Feature Refinement
While position information is encoded using the sinusoidal encoder described earlier, the image information is processed into high-dimensional features through a modified I3D model [23], in which the stride along the time axis is removed because of the limited frames. By concatenating the motion and image features together, a multi-modal feature is formed. The Temporal Feature Refinement module is designed to integrate these multi-modal features into a unified representation and upsample them to the time scale of the graph motion sequence.
As demonstrated in Fig. 1, the module begins by averaging the encoded position features to match the temporal-spatial resolution of the image features. Then the concatenated features are refined by three consecutive bottleneck layers, where the features are compressed in half at the bottleneck place and then remapped back to the high dimensions. The refined features are further upsampled to the motion time scale through a temporal pyramid pooling layer [12] with four parallel temporal pooling blocks. The final refined features are obtained by introducing the interpolated original features into the pooled features.
@@ III-C Multi-Modal Graph Convolutional Network
In addition to the motion and image feature refinement stream, the original motion sequence is processed in a graph encoder-decoder network [1]. This network models the structural dependencies between motion elements (skeleton and objects) by representing them as spatial and temporal graphs. The graph nodes are skeleton keypoints and object center points, and the graph edges represent the dynamic relations between nodes. Through multiple layers of graph convolutions and pooling operations, the network learns high-dimensional motion features that preserve both local and global motion contexts. The output of the graph encoder-decoder has the same dimensionality as the motion feature refinement stream, allowing for seamless fusion.
@@ V EXPERIMENTS AND RESULTS
To evaluate the performance of the proposed modules, we conduct the experiments on a public human-objects interaction dataset: Bimanual Actions [24]. On this dataset, we also perform systematic ablation studies to validate four critical design choices:
@@ V-A Datasets
Bimanual Actions dataset [24] captures fine-grained human-object interactions for kitchen tasks comprising 540 recordings (2 hours 18 minutes total) of bimanual manipulation tasks such as cutting and stirring. The dataset authors propose a leave-one-subject-out cross-validation protocol, where models are evaluated on held-out participants to test generalization to unseen users. For our experiments, we compare against state-of-the-art methods on this benchmark while conducting ablation studies on Subject 1’s testset to isolate the impact of our proposed components (e.g., label smoothing, fusion strategies).
@@
[…]
@@
    a ”Smoothing” means Label Smoothing, ”Mixing” represents Weighted Mixing, ”Sinusoidal” is the Sinusoidal Encoder, and ”Refinement” is the Temporal Feature Refinement block.
@@
    b ”O” is the original label, ”L” means the linear filtered label, and ”G” means the Gaussian filtered label. ”w” is with and ”w/o” means without.
@@ V-B Experimental Settings
We evaluate our method using standard metrics to ensure a comprehensive assessment of its performance. For evaluation, we employ metrics such as recognition accuracy, micro-F1, and segmentation accuracy (F1@k). The model is optimized using the SGD optimizer with an initial learning rate of $10^{-4}$, which is decayed following a multi-step schedule. Training is conducted on two NVIDIA RTX 2080ti GPUs, leveraging mixed precision to enhance efficiency and reduce memory overhead. We use a batch size of 32 and train for 60 epochs, applying the SmoothLabelMix data augmentation techniques to improve generalization. The evaluation experiments are conducted on a single GPU.
@@ V-C Ablation Study
Our ablation study systematically evaluates the contributions of three core innovations in the framework: SmoothLabelMix data augmentation, the Temporal Refinement module, and the Sinusoidal Encoder within the multi-modal graph convolutional network. The introduction of the image naturally raises the challenge of fusion strategies. Therefore, while analyzing the Sinusoidal Encoder, we also explored the impact of different fusion approaches: early, middle, late, and middle+late. Note that we categorize the implementation of the Sinusoidal Encoder and Temporal Refinement as a middle fusion method, because the Sinusoidal Encoder processes the motion features before fusion within the Temporal Refinement.
Beginning at the $7$-th row, we introduce the image stream alongside the motion sequence as multi-modal inputs. Rows 7-8 of the ablation study (Table I) evaluate the impact of the proposed Temporal Refinement module, which enhances the integration of visual and motion modalities. Integrating the Temporal Refinement module improves the frame recognition accuracy by $1.23\%$, while also mitigating over-segmentation, particularly beneficial for F1@50 ($+2.00\%$). The results demonstrate the module’s effectiveness in enhancing temporal coherence and segmentation precision. However, the inclusion of image streams still brings disturbance to the model, especially for action segmentation. This observation motivates us to further analyze the configuration of the fusion strategy and the Sinusoidal Encoder.
TABLE II: Comparison of cross-validation results with state-of-the-art methods on the Bimanual Actions dataset [24]a
[…]
Dreher et al. [24] $63.0$ $64.0$ $40.6\pm 7.2$ $34.8\pm 7.1$ $22.2\pm 5.7$
H2O+RGCN [25] 68.0 66.0 -- -- --
Independent BiRNN [26] $74.8$ $76.7$ $74.8\pm 7.0$ $72.0\pm 7.0$ $61.8\pm 7.3$
Relational BiRNN [26] $77.5$ $80.3$ $77.7\pm 3.9$ $75.0\pm 4.2$ $64.8\pm 5.3$
ASSIGN [26] $82.6$ $79.8$ $84.0\pm 2.0$ $81.2\pm{2.0}$ $68.5\pm{3.3}$
2G-GCN [27] -- -- $85.0\pm 2.2$ $82.0\pm 2.6$ $69.2\pm 3.1$
PGCN [12] 86.886.8{86.8}86.8 83.983.9{83.9}83.9 ${88.5}\pm{\bm{1.1}}$ ${85.5}\pm{2.0}$ ${77.0}\pm 3.4$
TFGCN [1] 88.488.4{88.4}88.4 88.688.6{88.6}88.6 ${93.7}\pm 1.2$ ${91.9}\pm{\mathbf{1.6}}$ ${85.4}\pm{\mathbf{2.9}}$
@@
Figure 3: Qualitative comparison of action segmentation results: Visualized segmentation outputs of different models for the breakfast cereal preparation scenario [24], illustrating the temporal alignment and accuracy of predicted action boundaries.

### v2

@@ I INTRODUCTION
Our preliminary studies [xing2024understanding] on human-robot collaboration reveal that encoder-decoder graph convolutional networks excel at parsing spatiotemporal information by representing human skeleton and object center points as graph nodes. Through attention mechanisms, these networks dynamically update the relationships between nodes, capturing the evolving interactions between human and objects. However, joint misdetections and tracking errors in skeleton data often lead to over-segmentation, fragmenting continuous actions into disjointed segments and disrupting the temporal coherence of action sequences. This limitation becomes particularly critical in human-object interaction scenarios, where both human pose estimation (skeleton) and object detections (bounding boxes) introduce noise. Object misdetections, such as missing and inaccurately localized bounding boxes, lead to inconsistencies in action segmentation.
@@
Recent advances in multi-modal networks [duan2022revisiting, tani2024graph] have demonstrated promise in addressing noise-induced over-segmentation by integrating structured 3D motion data (e.g., skeletal trajectories and object centroids) with pixel-based visual streams. However, integrating pixel-based RGB features with 3D position information presents significant challenges. The two modalities exhibit divergent characteristics: 3D position data are spatially sparse and structured, while image features are spatially dense and unstructured. Directly concatenating or aligning these representations leads to feature incompatibility. To bridge this gap, we introduce a sinusoidal joint encoding strategy that projects 3D coordinates into a continuous sin-cos space. This transformation downscales noisy joint trajectories while preserving spatial relationships, enabling seamless fusion with visual features.
@@
A second challenge arises from computational efficiency: processing full-frame video data at high temporal resolutions is prohibitively expensive for real-time systems, while naively fusing features across different temporal resolutions degrades performance. A common approach is to upsample low-frequency image features to match the motion sequence’s time scale and merge them using a late fusion strategy [tani2024graph]. However, this final fusion step often results in feature loss due to the mismatch in the fineness of features, and it overlooks fine-grained temporal dependencies and misaligns crucial visual and motion cues. To address this, we propose a Multi-Modal Graph Convolutional Network (MMGCN) that processes high-frequency motion data via a hierarchical graph convolutional network and strategically integrates high-frequency motion features into low-frequency (1 fps) visual cues at a middle stage, ensuring a more coherent fusion with processed motion information later in the pipeline. By processing the two modalities in different temporal resolutions, our method effectively preserves fine-grained temporal dependencies and aligns visual motion cues efficiently.
@@ II-A Human Action Segmentation
Human Action Segmentation, which aims to temporally localize and classify sub-actions within continuous sequences, has evolved from early template-based methods [patron2010high, xu2012streaming, xing2021robust] to modern image-based learning architectures [lea2017temporal, farha2019ms]. Temporal Convolutional Networks (TCNs) [lea2017temporal] pioneered frame-wise prediction using dilated convolutions, while MS-TCN [farha2019ms] extended this with multi-stage refinement. Recent works like Action Segment Refinement Framework [ishikawa2021alleviating] combine boundary detection and classification to improve temporal localization. Despite progress, these methods suffer from high computational costs, sensitivity to background clutter, and the inclusion of redundant visual information, making them inefficient for real-time applications.
To address these challenges, graph-based networks are considered as a more efficient alternative by modeling human motion as structured spatial and temporal graphs, significantly reducing background noise and computational overhead. Early works, such as Spatial-Temporal Graph Convolutional Network (ST-GCN) [yan2018spatial], introduced graph convolutional networks to sequentially capture joint-level motion patterns across both spatial and temporal dimensions. Subsequent improvements, like Two-Stream Adaptive Graph Convolutional Network (2s-AGCN) [shi2019two] and Multi-Scale Graph Network (MS-G3D) [liu2020disentangling], incorporated attention mechanisms and multi-scale feature extraction to enhance action recognition. More recently, methods such as Pyramid Graph Convolutional Network (PGCN) [xing2022understanding] and Temporal Fusion Graph Convolutional Network (TFGCN) [xing2024understanding] have explored temporal modeling techniques to further improve segmentation accuracy.
@@ II-B Multi-Modal Models
Multi-modal learning is a promising method for action recognition, leveraging complementary strengths of visual (RGB/depth) and skeletal modalities to overcome individual limitations. Early works, like [luo2014spatio, shahroudy2015multimodal] fuse images/depths and skeletons per frame to improve the recognition results. Recent advancements focus on intermediate fusion strategies to better align cross-modal features. Kim et al. [kim2023cross] proposed a cross-modal transformer that attends to spatial correspondences between skeleton joints and image regions, improving recognition consistency by dynamically weighting discriminative features. Similarly, Wang et al. [wang2024cmf] integrates skeleton joints with RGB patches using a late fusion technique, but its reliance on dense pixel processing limits scalability. To address efficiency, PoseC3D [duan2022revisiting] encodes skeletons as pseudo-3D heatmaps and fuses them with RGB features via 3D convolutions, achieving a balance between accuracy and speed. More recently, Tani Hiroaki [tani2024graph] introduced a temporal attention mechanism that selects one representative image for the entire motion sequence capturing the minimal appearance features necessary for action recognition. While this approach is effective for short-term action recognition, it becomes computationally intensive for long-duration actions.
@@ II-C Data Augmentation
Data augmentation plays a pivotal role in improving the generalization and robustness of temporal models, particularly for action recognition and segmentation tasks. Traditional techniques, such as spatial augmentations (e.g., cropping, flipping) [simonyan2014two] and temporal perturbations (e.g., frame skipping, time warping) [wang2016temporal], focus on enhancing spatial invariance or simulating temporal variations. While effective for classification, these methods often disrupt the temporal coherence required for precise segmentation, where smooth transitions between sub-actions are critical.
Recent approaches adapt image-domain augmentation strategies to temporal tasks. Mixup [zhang2018mixup] and CutMix [yun2019cutmix], which linearly interpolate inputs and labels, have been extended to videos by blending random clips (e.g., VideoMix [yun2020videomix]). However, these methods prioritize spatial or short-term temporal invariance, inadvertently introducing abrupt label transitions that exacerbate over-segmentation artifacts. For instance, TemporalMix [xiang2024joint] downsamples two segments by half, interpolates the larger segment to match the smaller one to preserve the natural smoothness of human actions, and then combines them into a new sequence. However, its downsampling process disrupts the natural tempo of human actions, leading to inconsistent predictions in speed-sensitive tasks.
@@ III Multi-Modal Graph Convolutional Network
The idea of the Multi-Modal Graph Convolutional Network is jointly optimizing dense spatial features from visual data and dense temporal dynamics from motion sequences (skeleton and object). Motion inputs (e.g., 30 fps) provide fine-grained motion patterns but suffer from joint tracking noise, while sparse RGB frames (1 fps) offer stable spatial context but lack temporal granularity. Our model bridges these modalities through two core mechanisms: a sinusoidal encoder and temporal feature refinement.
@@ III-A Sinusoidal Encoder
To mitigate spatial jitter and enhance compatibility with visual features, we transform the joint position information in the camera coordinate into a continuous sin-cons representation. Given a 3D joint $P_{i}=(X_{i},Y_{i},Z_{i})$, where $P_{i}\in\mathbb{R}^{1\times 3}$, the encoder maps each coordinate into a high-dimensional space using sinusoidal and cosine functions. The encoder first scales these coordinates using a frequency factor determined by the parameters $\alpha$ and $\beta$. Then it computes sinusoidal sin_esin\_e and cosine con_econ\_e embeddings for each dimension, resulting in a high-dimensional vector that captures both the position and its spatial context. Mathematically, for each coordinate $c\in\{X_{i},Y_{i},Z_{i}\}$, the encoder computes:
@@
We set $d=128$ and obtain a $768$ dimensional embedded feature map ($3$ coordinates $\times 128\times 2$ functions). The $\alpha$ is set as $1000$ to create geometrically increasing wavelengths from $1$ to $1000$ units, capturing both fine and coarse-grained spatial features. The $\beta=100$ scales the values of the real-world coordinates. A $1\times 1$ convolution then compresses this to a target dimension ($512$ channels), balancing expressiveness and efficiency.
While transformer-based models like GCNs employ learned positional encodings to adaptively update inter-node relations through attention mechanisms, this flexibility comes with significant computational costs. In contrast, the sinusoidal encoder provides a fixed geometric prior that explicitly preserves spatial relations through wavelength-scaled trigonometric function. The fixed encoding prevents overfitting to sparse training skeletons while maintaining spatial continuity between positions. This proves particularly valuable when aligning with sparse visual features where spatial precision matters more than temporal sequencing. More evidence can be found in the ablation studies in the Section V.
@@ III-B Temporal Feature Refinement
While position information is encoded using the sinusoidal encoder described earlier, the image information is processed into high-dimensional features through a modified I3D model [carreira2017quo], in which the stride along the time axis is removed because of the limited frames. By concatenating the motion and image features together, a multi-modal feature is formed. The Temporal Feature Refinement module is designed to integrate these multi-modal features into a unified representation and upsample them to the time scale of the graph motion sequence.
As demonstrated in Fig. 1, the module begins by averaging the encoded position features to match the temporal-spatial resolution of the image features. Then the concatenated features are refined by three consecutive bottleneck layers, where the features are compressed in half at the bottleneck place and then remapped back to the high dimensions. The refined features are further upsampled to the motion time scale through a temporal pyramid pooling layer [xing2022understanding] with four parallel temporal pooling blocks. The final refined features are obtained by introducing the interpolated original features into the pooled features.
@@ III-C Multi-Modal Graph Convolutional Network
In addition to the motion and image feature refinement stream, the original motion sequence is processed in a graph encoder-decoder network [xing2024understanding]. This network models the structural dependencies between motion elements (skeleton and objects) by representing them as spatial and temporal graphs. The graph nodes are skeleton keypoints and object center points, and the graph edges represent the dynamic relations between nodes. Through multiple layers of graph convolutions and pooling operations, the network learns high-dimensional motion features that preserve both local and global motion contexts. The output of the graph encoder-decoder has the same dimensionality as the motion feature refinement stream, allowing for seamless fusion.
@@ V EXPERIMENTS AND RESULTS
To evaluate the performance of the proposed modules, we conduct the experiments on a public human-objects interaction dataset: Bimanual Actions [dreher2019learning]. On this dataset, we also perform systematic ablation studies to validate four critical design choices:
@@ V-A Datasets
Smoothingb Mixing Refinement Sinusoidal Fusion Strategies c Evaluation Metrics ($\%$) d
@@
    a “Smoothing” means Label Smoothing, “Mixing” represents Weighted Mixing, “Sinusoidal” is the Sinusoidal Encoder, and “Refinement” is the Temporal Feature Refinement block.
@@
    b “O” is the original label, “L” means the linear filtered label, and “G” means the Gaussian filtered label. “w” is with and “w/o” means without.
@@
Bimanual Actions dataset [dreher2019learning] captures fine-grained human-object interactions for kitchen tasks comprising 540 recordings (2 hours 18 minutes total) of bimanual manipulation tasks such as cutting and stirring. This dataset provides comprehensive framewise annotations including 3D bounding boxes for 12 objects, 3D skeletons for 6 human subjects and 14 interaction categories (including both-hand interactions). For our experiments, we represent each 3D skeleton using 15 upper-body joints (5 head joints, 6 arm joints, and 4 torso/hip joints) and represent objects by their 3D bounding box centroids. Following the dataset authors’ protocol, we employ leave-one-subject-out cross-validation to evaluate generalization to unseen users. For ablation studies, we specifically evaluate on Subject 1’s test set to isolate the impact of proposed components (e.g., label smoothing, fusion strategies), while comparing against state-of-the-art methods on the full benchmark.
@@ V-B Experimental Settings
We evaluate our method using standard metrics to ensure a comprehensive assessment of its performance. For evaluation, we employ metrics such as recognition accuracy, micro-F1, and segmentation accuracy (F1@k). The model is optimized using the SGD optimizer with an initial learning rate of $10^{-4}$, weight decay of $0.0005$, and multi-step learning rate decay at epochs 20 and 40. Training runs for 60 epochs on two NVIDIA RTX 2080ti GPUs, leveraging mixed precision to enhance efficiency and reduce memory overhead. We apply the SmoothLabelMix data augmentation techniques to improve generalization. The evaluation experiments are conducted on a single GPU.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 3: Three fusion strategies with the Temporal Feature Refinement module: (a) early fusion, (b) middle fusion, (c) later fusion. I: image sequence, M: motion sequence, IE: image encoder, TFR: temporal feature refinement, GE: graph encoder, GD: graph decoder, CF: classifier, P: predictions.
@@ V-C Ablation Study
Our ablation study systematically evaluates the contributions of three core innovations in the framework: SmoothLabelMix data augmentation, the Temporal Refinement module, and the Sinusoidal Encoder within the multi-modal graph convolutional network. The introduction of the image naturally raises the challenge of fusion strategies. Therefore, while analyzing the Sinusoidal Encoder, we also explored the impact of different fusion approaches: early, middle, late, and middle+late. Figure 3 illustrates three representative examples of these strategies integrated with the Temporal Feature Refinement block. These strategies differ primarily in the stage at which image and motion features are fused into the main processing stream. For configurations omitting the Temporal Feature Refinement module, the encoded features are upsampled via a linear interpolation layer instead. Note that we categorize the implementation of the Sinusoidal Encoder as a middle fusion method, because the Sinusoidal Encoder processes the motion features before fusion within the Temporal Refinement.
TABLE II: Comparison of cross-validation results with state-of-the-art methods on the Bimanual Actions dataset [dreher2019learning]a
Model Accuracy ($\%$) F1 macro ($\%$) F1@10 ($\%$) F1@25 ($\%$) F1@50 ($\%$)
Dreher et al. [dreher2019learning] $63.0$ $64.0$ $40.6\pm 7.2$ $34.8\pm 7.1$ $22.2\pm 5.7$
H2O+RGCN [lagamtzis2023exploiting] 68.0 66.0 - - -
Independent BiRNN [morais2021learning] $74.8$ $76.7$ $74.8\pm 7.0$ $72.0\pm 7.0$ $61.8\pm 7.3$
Relational BiRNN [morais2021learning] $77.5$ $80.3$ $77.7\pm 3.9$ $75.0\pm 4.2$ $64.8\pm 5.3$
ASSIGN [morais2021learning] $82.6$ $79.8$ $84.0\pm 2.0$ $81.2\pm{2.0}$ $68.5\pm{3.3}$
2G-GCN [qiao2022geometric] - - $85.0\pm 2.2$ $82.0\pm 2.6$ $69.2\pm 3.1$
PGCN [xing2022understanding] 86.8{86.8} 83.9{83.9} ${88.5}\pm{\bm{1.1}}$ ${85.5}\pm{2.0}$ ${77.0}\pm 3.4$
TFGCN [xing2024understanding] 88.4{88.4} 88.6{88.6} ${93.7}\pm 1.2$ ${91.9}\pm{\mathbf{1.6}}$ ${85.4}\pm{\mathbf{2.9}}$
@@
Beginning at the $7$-th row, we introduce the image stream alongside the motion sequence as multi-modal inputs. Rows 7-8 of the ablation study (Table I) evaluate the impact of the proposed Temporal Refinement module, which enhances the integration of visual and motion modalities. Integrating the Temporal Refinement module improves the frame recognition accuracy by $1.23\%$, while also mitigating over-segmentation, particularly beneficial for F1@50 ($2.00\%$). The results demonstrate the module’s effectiveness in enhancing temporal coherence and segmentation precision. However, the inclusion of image streams still brings disturbance to the model, especially for action segmentation. This observation motivates us to further analyze the configuration of the fusion strategy and the Sinusoidal Encoder.
The middle, and late fusion variants all fuse sparse appearance features with motion representations processed through a transformer-based encoder (Graph Encoder). The middle+late fusion substitutes this transformer encoder with our Sinusoidal Encoder, which operates as a middle fusion component by processing raw coordinates before temporal refinement. This architectural distinction proves that the Sinusoidal Encoder’s fixed geometric priors outperform learned transformer embeddings in noise suppression while maintaining spatial coherence. It enables a more effective alignment with sparse visual features.
For efficiency analysis, our model archives an optimal balance between computational efficiency and performance, operating at $131.0$ GFLOPs, significantly lower than using higher visual temporal resolution. For instance: processing motion and visual modalities at a resolution ratio of $30:2$ increases FLOPs to $220.0$ G, while uniform high-resolution processing ($30:30$) escalates costs to $2687.3$ GFLOPs, a $20.5\times$ increase. This exponential scaling highlights the inefficiency of naively aligning modalities at high frequencies. While our framework introduces moderate computational overhead compared to models using pure motion sequences, it retains real-time performance (81.0 FPS on a single GPU) by strategically decoupling temporal resolutions.
V-D Comparison with states-of-the-art
Figure 4: Qualitative comparison of action segmentation results: Visualized action segmentation outputs of different models for the left arm in the breakfast cereal preparation scenario [dreher2019learning], illustrating the temporal alignment and accuracy of predicted action boundaries. Top: input frames with perception artifacts (a-d). Bottom: ground-truth and predictions.